OpenStudy (acespeedfighter):

Which of the following is the solution to the equation x + 8.6 = 36.2? (3 points) 10.4 17.2 27.6 44.8

rebeccaxhawaii (rebeccaxhawaii):

Get x by itself so subtract 8.6 from both sides.

OpenStudy (kittiwitti1):

\[x+8.6=36.2\]\[x+8.6-8.6=36.2-8.6\]\[x\cancel{+8.6}\cancel{-8.6}=36.2-8.6\]\[x+0=36.2-8.6\]\[x=36.2-8.6\]

OpenStudy (acespeedfighter):

27.6 is the answer right?

OpenStudy (kittiwitti1):

\(\text{Yes, good job}\) ☺

OpenStudy (acespeedfighter):

thx :)